OpenStudy (anonymous):

please help! ln(x/2)+ln(x+3/5)=0 i got up to ln[(x^2+3x)/10]=0

OpenStudy (anonymous):

that makes \[\frac{x^2+3x}{10}=1\]

OpenStudy (anonymous):

since \(\ln(\xi)=0\iff \xi=1\)

OpenStudy (anonymous):

do you multiply the 10 by both sides so you have x^2+3x=10

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

idk how to figure it out